Use the definition of big-O to prove that 12 + 22 + ... + n2 is O(n3)
Answer #1

1^2+2^2+3^2+.....+n^2=((n*(n+1)*(2*n+1)))/6

                                          =((n^2+n)*(2*n+1))/6

                                          =(2*n^3+n^2+2*n^2+n)/6

                                          =(2*n^3+3*n^2+n)/6

f(n)=(2*n^3+3*n^2+n)/6

g(n)=n^3

c=3

n>=n0 = 1

(2*n^3+3*n^2+n)/6 <= 3*n^3

So time complexity of 1^2+2^2+3^2+.....+n^2 is O(n^3)